Michelle is eight years older than Adam. Three years ago Michelle was twice as old as Adam was then. How old is Michelle?
Answer by Earlsdon(6294)

Let M = Michelle's present age.
Let A = Adam's present age.
M = A + 8 Michelle is 8 yrs older than Adam.
M-3 = 2(A-3) 3 yrs ago, Michelle was twice Adam's age. Solve this equation for A and substitute into the first equation, then solve for M.
M-3 = 2A-6 Add 6 to both sides.
M+3 = 2A Divide both sides by 2
M/2+3/2 = A Substitute this into the first equation and solve for M
M = ((M/2)+3/2)+8 Multiply by 2 to clear the fractions.
2M = (M+3) + 16 Simplify.
M = 19
Michelle's age is 19.